# Are Statistical Methods Obsolete in the Era of Deep Learning?

### v1

@@ 5 Discussion
Figure S4: Reconstructed trajectories by PINN and MAGI on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. Each solid blue curve is the estimate from one data set. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
@@
Figure S5: Forecasted trajectories by PINN and MAGI on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be predicted. Each solid blue curve is the prediction from one data set. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: MAGI estimates. Lower five rows: PINN estimates over all five tested $\lambda$ hyperparameter values.
S5.1 Results from stable (transient chaos) regime
@@
Our second testbed, depicted in SI Figure S6, which we refer to as “Stable (Transient Chaos),” initially shows butterfly oscillatory behavior, before eventually converging to a fixed point. This regime is mathematically stable (Sparrow, 2012), with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},23,10)$ and initial condition $(X(0),Y(0),Z(0))=(5,5,5)$.
@@
Mirroring the “Chaotic (Butterfly)” testbed, for “Stable (Transient Chaos),” we generate 100 independent data sets. The observation window is from $t\in[0,8]$, and a total of 81 equally-spaced observations are generated within this interval for each data set, which gives $I_{\text{obs}}$, with $5\%$ additive Gaussian noise injected. For both PINN and MAGI, we use a shared discretization set of $I$ as $321$ evenly-spaced time steps in $t\in[0,8]$ to evaluate the physics-based loss component (see Equation (3)) and $W_{I}$ (see Equation (4), respectively.
In the Stable (Transient Chaos) regime, depicted in SI Figure S6, we observe similar behaviors in MAGI and PINN as in the Chaotic (Butterfly) regime. The flat-line output of PINN remains evident across a wide range of $\lambda$ values, as shown in SI Figure S7, although it occurs less frequently compared to the chaotic regime. MAGI continues to perform consistently well across datasets and time periods.
@@
A closer analysis of trajectory RMSE, presented in the boxplots of SI Figure S8, confirms this visual impression. MAGI achieves the smallest error, while PINN at $\lambda=10$ shows an error closer to MAGI. The outlier error metrics in the boxplot correspond to instances of underfitted flat-line outputs, but the majority of PINN at $\lambda=10$ results across 100 datasets now exhibit reasonable performance. The physics-emphasizing PINN variants with small $\lambda$ continue to produce trivial flat-line predictions, which significantly contribute to the large errors observed in the boxplots.
@@
Figure S7: MAGI and PINN reconstructed trajectories on the Lorenz Stable (Transient Chaos) testbed across $100$ datasets, with one example dataset visualized. The dashed black lines give the true curves, which are to be inferred. Each solid blue curve is the estimate from one data set. Top row: the $100$ MAGI posterior means for each dataset. Bottom two rows: the $100$ PINN forward-pass outputs for each dataset, under hyperparameter settings of $\lambda=10$ and $\lambda=1000$. We note that the 100 reconstructed trajectories for MAGI are all indeed in the subplots: they are all on top of each other, indicating very low variance across datasets.
@@
Figure S8: Boxplots showing trajectory reconstruction RMSEs across $100$ datasets on the Lorenz Stable (Transient Chaos) testbed. Lower values indicate better performance. The three panels correspond to the three system components $X,Y$, and $Z$. In each panel, the leftmost blue boxplot is for MAGI, while the remaining red boxplots are for PINN under different hyperparameter settings. The dashed vertical line separates the MAGI and PINN results.
For parameter inference, SI Figure S9 presents the boxplots of absolute errors for each model parameter. As expected, MAGI continues to demonstrate better accuracy in parameter estimation. However, the performance of PINN on $\sigma$ and $\rho$ has improved compared to the chaotic regime, likely due to enhanced accuracy in trajectory recovery under the stable regime.
Although the flat-line behavior of PINN can be explained, the causes of the improved performance of PINN in the Stable (Transient Chaos) regime compared to the chaotic regime remain unclear, largely due to the well-known challenges in interpreting neural networks. This lack of interpretability leaves the underlying reasons for this behavior open to speculation.
@@
Figure S9: Boxplots showing parameter inference absolute errors across $100$ datasets on the Lorenz Stable (Transient Chaos) testbed. Lower values indicate better performance. The three panels correspond to the three system parameters $\beta,\rho$, and $\sigma$. In each panel, the leftmost blue boxplot is for MAGI, while the remaining red boxplots are for PINN under different hyperparameter settings. The dashed vertical line separates the MAGI and PINN results.
@@
The future prediction results in SI Figure S10 closely resemble those of the chaotic (butterfly) regime. All PINN variants still exhibit a tendency to drift toward flat-line predictions as the forecast extends further into the future. In contrast, MAGI demonstrates consistent and accurate forecasting over the entire interval from $t=2$ to $t=5$. However, the forecasts begin to diverge noticeably near the end of the interval at $t=5$, reaching the limit of this Lorenz system example.
@@
Figure S10: MAGI and PINN forecasted trajectories on the Stable (Transient Chaos) testbed across $100$ datasets, with one example dataset visualized. The dashed black lines give the true curves, which are to be predicted. Each solid blue curve is the prediction from one data set. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: the $100$ MAGI posterior means for each dataset. Bottom two rows: the $100$ PINN forward-pass outputs for each dataset, under hyperparameter settings of $\lambda=10$ and $\lambda=1000$.
@@
Comparing the results between the two testbeds, one may hypothesize that forecasting is more tractable on the Chaotic (Butterfly) regime than the Stable (Transient Chaos) regime, the former which we analyzed at the end of Section 4. Overall, from the above results, it is clear that the Chaotic (Butterfly) and Stable (Transient Chaos) testbeds tell very similar stories about the performance differences between MAGI and PINN.

### v2

@@ 5 Discussion and Conclusion
Figure S8: Reconstructed trajectories by best-case-per-$\lambda$ PINN and MAGI on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be identified. Each solid blue curve is the estimate from one data set. Top row: MAGI estimates. Lower five rows: best-case-per-$\lambda$ PINN estimates over all five tested $\lambda$ hyperparameter values.
@@
Figure S9: Forecasted trajectories by best-case-per-$\lambda$ PINN and MAGI on the Lorenz Chaotic (Butterfly) testbed across $100$ datasets, with one example dataset visualized. The dots show one sample dataset (out of 100). The dashed black lines give the true curves, which are to be predicted. Each solid blue curve is the prediction from one data set. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: MAGI estimates. Lower five rows: best-case-per-$\lambda$ PINN estimates over all five tested $\lambda$ hyperparameter values.
Figure S10: PINN results from the practitioner-style implementation (van Herten et al. 2022) for the Lorenz Chaotic (Butterfly) testbed, which performs worse than the DeepXDE-based PINN used in the main text. Boxplots showing trajectory reconstruction RMSEs across $100$ datasets. Lower values indicate better performance. The three panels correspond to the system components $X$, $Y$, and $Z$. The legend and panel layout are identical to Figure 9 in the main text; see the caption there for details. In each panel, the leftmost boxplot corresponds to MAGI, and the remaining boxplots correspond to practitioner-style PINN under different hyperparameter values $\lambda$. The dashed vertical line separates MAGI and PINN results.
Figure S11: PINN results from the practitioner-style implementation (van Herten et al. 2022) for parameter estimation on the Lorenz Chaotic (Butterfly) testbed, which yields worse performance than the DeepXDE-based PINN used in the main text. Boxplots showing parameter inference absolute errors across $100$ datasets. Lower values indicate better performance. The three panels correspond to the system parameters $\beta$, $\rho$, and $\sigma$. The legend and panel layout are identical to Figure 10 in the main text; see the caption there for details. In each panel, the leftmost boxplot corresponds to MAGI, and the remaining boxplots correspond to practitioner-style PINN under different hyperparameter values $\lambda$. The dashed vertical line separates MAGI and PINN results.
S6 Results from stable (transient chaos) regime
@@
Our second testbed, depicted in SI Figure S12, which we refer to as “Stable (Transient Chaos),” initially shows butterfly oscillatory behavior, before eventually converging to a fixed point. This regime is mathematically stable (Sparrow 2012), with parameters $\bm{\theta}=(\beta,\rho,\sigma)=(\frac{8}{3},23,10)$ and initial condition $(X(0),Y(0),Z(0))=(5,5,5)$.
@@
Mirroring the “Chaotic (Butterfly)” testbed, for “Stable (Transient Chaos),” we generate 100 independent data sets. The observation window is from $t\in[0,8]$, and a total of 81 equally-spaced observations are generated within this interval for each data set, which gives $I_{\text{obs}}$, with $5\%$ additive Gaussian noise injected. For both PINN and MAGI, we use a shared discretization set of $I$ as $321$ evenly-spaced time steps in $t\in[0,8]$ to evaluate the physics-based loss component (see Equation (3)) and $W_{I}$ (see Equation (4), respectively. All PINN and MAGI implementation details are identical to those described in Section 4.1.
In the Stable (Transient Chaos) regime, as shown in SI Figure S12, MAGI’s inferred trajectories continue to closely hug the ground-truth trajectory, with low trial-to-trial variability. In contrast, even the best-performing PINNs’ inferred trajectories appear noticeably “fuzzier,” with subtantially larger variability over trials. We also observe flat-line trajectory collapse for PINNs with smaller $\lambda$ values. In general, MAGI and PINN behaviors are very similar to those in in the Chaotic (Butterfly) regime.
@@
A closer analysis of trajectory RMSE, presented in the boxplots of SI Figure S14, confirms this visual impression. MAGI achieves the smallest error, while PINN at $\lambda=10$ shows an error closer to MAGI. The physics-emphasizing PINN variants with small $\lambda$ continue to produce trivial flat-line predictions, which significantly contribute to the large errors observed in the boxplots.
@@
Figure S13: MAGI and PINN (best case per $\lambda$) reconstructed trajectories on the Lorenz Stable (Transient Chaos) testbed across $100$ datasets, with one example dataset visualized. The dashed black lines give the true curves, which are to be inferred. Each solid blue curve is the estimate from one data set. Top row: the $100$ MAGI posterior means for each dataset. Lower five rows: best-case-per-$\lambda$ PINN estimates over all five tested $\lambda$ hyperparameter values. We note that the 100 reconstructed trajectories for MAGI are all indeed in the subplots: they are all on top of each other, indicating very low variance across datasets.
@@
Figure S14: Boxplots showing trajectory reconstruction RMSEs across $100$ datasets on the Lorenz Stable (Transient Chaos) testbed. Lower values indicate better performance. The three panels correspond to the three system components $X,Y$, and $Z$. In each panel, the leftmost blue boxplot is for MAGI, while the remaining red boxplots are for PINN (best case per $\lambda$) under different $\lambda$ hyperparameter settings. The dashed vertical line separates the MAGI and PINN results.
For parameter inference, SI Figure S15 presents the boxplots of absolute errors for each model parameter. Results are very similar to those of the Chaotic (Butterfly) setting: MAGI outperforms or performs comparably to the strongest PINN variants on $\beta$ and $\rho$, but underperforms slightly on $\sigma$.
@@
Figure S15: Boxplots showing parameter inference absolute errors across $100$ datasets on the Lorenz Stable (Transient Chaos) testbed. Lower values indicate better performance. The three panels correspond to the three system parameters $\beta,\rho$, and $\sigma$. In each panel, the leftmost blue boxplot is for MAGI, while the remaining red boxplots are for (best case per $\lambda$) under different $\lambda$ hyperparameter settings. The dashed vertical line separates the MAGI and PINN results.
@@
The future prediction results in SI Figure S16 are also similar to those of the Chaotic (Butterfly) regime. Here, all PINN variants exhibit a tendency to collapse towards flat-line predictions as the forecast extends further into the future. In contrast, MAGI demonstrates consistent and accurate forecasting over the entire interval from $t=2$ to $t=5$. However, the forecasts begin to diverge noticeably near the end of the interval at $t=5$, reaching the limit of this Lorenz system example. Comparing the results between the two testbeds, one may hypothesize that forecasting is more tractable on the Chaotic (Butterfly) regime than the Stable (Transient Chaos) regime, the former which we analyzed at the end of Section 4.
@@
Figure S16: MAGI and PINN (best case per $\lambda$) forecasted trajectories on the Stable (Transient Chaos) testbed across $100$ datasets, with one example dataset visualized. The dashed black lines give the true curves, which are to be predicted. Each solid blue curve is the prediction from one data set. The red dashed vertical line separates the in-sample observation period from the future forecasting period. Top row: the $100$ MAGI posterior means for each dataset. Lower five rows: best-case-per-$\lambda$ PINN estimates over all five tested $\lambda$ hyperparameter values.
@@
Overall, from the above results, it is clear that the Chaotic (Butterfly) and Stable (Transient Chaos) testbeds tell very similar stories about the performance differences between MAGI and PINN.